OpenStudy (anonymous):

fan and medal Which line is parallel to the line that passes through the points (1, 7) and (-3, 4)? y= -3/4x -5 y= 3/4x+1 y=4/3x-8 y=11/4+3

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

first we need to find the slope because a parallel line will have the same slope. We will use the slope formula : slope = (y2 - y1) / (x2 - x1) (1,7)......x1 = 1 and y1 = 7 (-3,4)...x2 = -3 and y2 = 4 now we sub into the slope formula... slope(m) = (4 - 7) / (-3-1) slope(m) = -3/-4 slope(m) = 3/4 now we will use y = mx + b slope(m) = 3/4 you can use either of your points...(-3,4)...x = -3 and y = 4 and since we already know our slope and we have our points (x,y)...we just need to find b, the y intercept. now we sub.. y = mx + b 4 = 3/4(-3) + b 4 = -9/4 + b 4 + 9/4 = b 16/4 + 9/4 = b 25/4 = b so our parallel equation is : y = 3/4x + 25/4 hmm....not an answer choice. I can tell you that parallel lines will have the same slope and the only answer choice tat has the same slope is : y = 3/4x + 1...so I would go with that one.

OpenStudy (anonymous):

okay thankyou ill hopeforthe best

OpenStudy (texaschic101):

sure thing :) good luck